3) Prove by induction that for all natural numbers n:


(f) n3 + 2n is divisible by 3.



1
Expert's answer
2021-05-31T19:36:06-0400

Let us prove by induction that "n^3 + 2n" is divisible by 3 for all natural numbers "n".


For "n=1" we have that "1^3+2\cdot 1=3" is divisible by 3.


Suppose that for "n=k" we have that "k^3 + 2k" is divisible by 3.


Let us prove the statement for "n=k+1."

Taking into account that

"(k+1)^3+2(k+1)=k^3+3k^2+3k+1+2k+2=(k^3+2k)+3(k^2+k+1)"

and "k^3+2k" and "3(k^2+k+1)" are divisible by 3, we conclude that "(k+1)^3+2(k+1)" is also divisible by 3.


We conclude that "n^3 + 2n" is divisible by 3 for all natural numbers "n".